OpenStudy (omgthatgirl):

Use the quadratic formula to solve the equation -2x^2+9x+3=0 a. 9+sqrt105/4 b. 9+sqrt52/2 c. 9+sqrt105/2 d. 4+sqrt210/18 <--- am i right

OpenStudy (mathteacher1729):

No, that's not correct. What did you use as your values of a, b, and c?

OpenStudy (omgthatgirl):

a=2 b=9 c=3

OpenStudy (mathteacher1729):

In the problem statement you wrote that a = -2. So I found the values for a = -2.

OpenStudy (omgthatgirl):

ok..

OpenStudy (omgthatgirl):

yea it is -2

OpenStudy (omgthatgirl):

@mathteacher1729

OpenStudy (mathteacher1729):

Ok, so you'll have to re-do your work and check again, this time let a = -2. :)

OpenStudy (mathteacher1729):

I have to get dinner now, good luck, I will try to sign on later to check in.

OpenStudy (omgthatgirl):

k thanks

OpenStudy (jdoe0001):

\(\bf \textit{quadratic formula}\\ {\color{blue}{ -2}}x^2{\color{red}{ +9}}x{\color{green}{ +3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\) solve for "x"

OpenStudy (omgthatgirl):

yea i did that and i got 19 for the denominator

OpenStudy (omgthatgirl):

i mean 18 sorry

OpenStudy (omgthatgirl):

and the numerator 81-24=57

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

19 for the denominator?

OpenStudy (jdoe0001):

how.. did you get 19 anyway for the denominator?

OpenStudy (omgthatgirl):

i meant 18

OpenStudy (jdoe0001):

well.. what's "a"?

OpenStudy (omgthatgirl):

-2

OpenStudy (jdoe0001):

so... "2a" would be?

OpenStudy (omgthatgirl):

oh i wasnt thinking. i thought the denominator was 2b my bad

OpenStudy (omgthatgirl):

2X-2=-4

OpenStudy (jdoe0001):

yeap thus \(\bf \textit{quadratic formula}\\ {\color{blue}{ -2}}x^2{\color{red}{ +9}}x{\color{green}{ +3}}=0 \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}} \\ \quad \\ x= \cfrac{ - {\color{red}{ 9}} \pm \sqrt { {\color{red}{ 9}}^2 -4{\color{blue}{ (-2)}}{\color{green}{ (3)}}}}{2{\color{blue}{ (-2)}}}\implies x=\cfrac{-9\pm\sqrt{81+24}}{-4}\)

OpenStudy (jdoe0001):

so... what would you end up in the numerator?

OpenStudy (omgthatgirl):

ok the answer is a

OpenStudy (jdoe0001):

dunno... what did you get on the numerator?

OpenStudy (omgthatgirl):

105 as the sqrt

OpenStudy (jdoe0001):

yeap thus \(\bf x= \cfrac{ - {\color{red}{ 9}} \pm \sqrt { {\color{red}{ 9}}^2 -4{\color{blue}{ (-2)}}{\color{green}{ (3)}}}}{2{\color{blue}{ (-2)}}}\implies x=\cfrac{-9\pm\sqrt{81+24}}{-4} \\ \quad \\ x=\cfrac{-9\pm\sqrt{105}}{-4}\implies x=\cfrac{9\mp\sqrt{105}}{4} \)

OpenStudy (omgthatgirl):

thanks

OpenStudy (jdoe0001):

yw